### v1

@@ 10 Conclusions
Quantum service orientation allows QC vendors to offer end-users’ access to computational resources - enabling shots at QPUs - by following the utility computing model. Recently, research and development on QSE have started to synergise the principles of service-orientation and practices of QC (algorithms, simulations, QuGates, etc.) to enable the adoption of QCaaS by individuals and enterprises to attain quantum supremacy in modern day computing. To this end, we used the systematic mapping study approach to primarily investigate (i) existing solutions for quantum service development, such as QSRs, modelling notations, patterns, and programming frameworks etc. to architect QCaaS, and (ii) emerging trends that highlight the needs for ongoing and future research on QCaaS. The results of this SMS as structured information (Table 2) and visual illustrations (Figure 2, 5 - Figure 7) highlight the strengths, limitations, and future research from QSE perspective.
@@
Needs for future work: The primary implication of this study is to establish an evidence-based body of knowledge for service development that leverages design notations, patterns, and architectural approaches highlighting the needs for human-centric and process-driven QCaaS development. The results of this SMS establish the foundations for future work in three directions that include (a) conducting practitioners’ survey, (b) implementing the reference architecture, and (c) mining social coding platforms for empiricism in QCaaS computing. The literary foundations can help us to design an empirical study that aims to engage service developers and engineers in a workshop (activity and survey) to seek their feedback and synthesise the results as practitioners’ perspectives to complement the evidence from academic research. We also plan to mine social coding platforms (e.g., GitHub) to empirically discover knowledge and understand the practices adopted by developers’ communities in open-source QCaaS.

### v2

@@ 10 Conclusions
Quantum service-orientation enables QC vendors to provide end-users with access to computational resources—allowing ‘shots’ on QPUs—through a utility computing model. Recent QSE research is combining service-orientation principles with quantum computing practices (algorithms, simulations, QuGates, etc.) to support QCaaS adoption by individuals and organisations, aiming for the anticipated leap into quantum computational supremacy. Using a systematic mapping study, we investigated (i) existing solutions for quantum service development, such as QSRs, modeling notations, patterns, and programming frameworks for architecting QCaaS, and (ii) emerging trends that highlight ongoing and future research needs. The results of this SMS as structured information (Table 2) and visual illustrations (Figure 2, 5 - Figure 7) highlight the strengths, limitations, and future research on QSE.
@@
Needs for future work: The results of this SMS provide a foundation for future work along three main directions: (a) conducting a practitioners’ survey, (b) validating the reference architecture based on different case studies, and (c) mining social coding platforms to strengthen empirical insights into QCaaS development. These findings can provided empirically-grounded foundations, employing mixed-method research approaches, thereby enriching the evidence base established through academic research and experimental analysis. Additionally, mining social coding platforms and open-source repositories (e.g., GitHub) will allow us to identify patterns, artifacts, and practices emerging from QCaaS developer communities.
Declaration of AI Assistance
During the preparation of this article, the authors used AI assistance to refine the language, perform grammar checking, and improve sentence structure. After the AI assistance, the authors proof-read, edited, and thoroughly reviewed the content of the article, wherever required. The authors take responsibility for the contents presented in this article.